OpenStudy (anonymous):

Solve: 112=-16t^2+128t

OpenStudy (anonymous):

I know you get two separate answers but I'm not sure the steps to get there.

hartnn (hartnn):

can you bring that in the form of \(ax^2+bx+c = 0 \)

hartnn (hartnn):

by adding 16t^2-128t on both sides.

OpenStudy (anonymous):

yes so, -16t^2+128t-112=0 or -16t^2-128t+112

OpenStudy (anonymous):

sorry i forgot the =0 on the second one

hartnn (hartnn):

ok, you brought 112 on other side, -16t^2 +128t-112 =0 is correct, but how about 16t^2 -128t+112 = 0 ?

OpenStudy (anonymous):

wouldn't you get the same answer?

hartnn (hartnn):

ofcourse ! and now can you simplify that a bit before proceeding ? notice that you can divide the entire equation by 16 and make those large numbers like 128,112....small

hartnn (hartnn):

16t^2 -16*8t +16*7 =0 ----->t^2 -8t +7=0 got this ?

OpenStudy (anonymous):

yes that's what i got

hartnn (hartnn):

t^2 -8t +7=0 is relatively simple to solve can you think of 2 numbers whose product is 7 and sum is -8 ??

OpenStudy (anonymous):

that would factor out to (t-7) (t-1) right?

hartnn (hartnn):

absolutely correct :)

hartnn (hartnn):

(t-7) (t-1) =0 gives you which 2 values of t ?

OpenStudy (anonymous):

1 and 7?

hartnn (hartnn):

good! thats your answer :)

OpenStudy (anonymous):

THANK YOU!

hartnn (hartnn):

welcome ^_^